# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ G.3 Bit complexity analysis of Block-Krylov PCA
In particular, we start with some desired bound $\epsilon^{\prime}_{0}\in(0,1)$ for $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|$ by setting the number of bits to $O(\log(\tfrac{qkmn\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}_{0}}))$. At each iteration we repeat Steps 1, 3, and 4 (we do not need to compute again the condition number of $\boldsymbol{\mathrm{K}}$ in Step 2). After at most $m=O(\log(1/\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})))$ iterations, the error $\epsilon^{\prime}_{m}$ satisfies $\epsilon^{\prime}_{m}=\Theta(\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}}))$, and we obtain two quantities $\widetilde{\operatorname{gap}}_{k}\in(1\pm\tfrac{1}{8})\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})$ and $\widetilde{\mu}_{k}\in\mu_{k}(\boldsymbol{\mathrm{M}})+\tfrac{1}{8}\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})$ with high probability. In each iteration we need to construct $\widetilde{\boldsymbol{\mathrm{M}}}$ using Steps 1, 3, and 4 with the specified number of bits.
Step 1 costs $O(qmnk^{\omega-2})$ operations. Step 3 requires $O(m(qk)^{\omega-1})$ operations. Step 4 executes $O(mn(qk)^{\omega-2})$ operations. The result is the $qk\times qk$ matrix $\widetilde{\boldsymbol{\mathrm{M}}}$ on which we call $\mathsf{COUNT}$ on each iteration. From Lemma E.1, assuming that we have regularized $\widetilde{\boldsymbol{\mathrm{M}}}$ appropriately using $\mathsf{REGULARIZE}$, $\mathsf{COUNT}$ costs at most $O((qk)^{\omega}\operatorname{polylog}(qk/\epsilon^{\prime}_{m}))=O((qk)^{\omega}\operatorname{polylog}(qk/\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})))$. The maximum number of bits is in the last iteration, which is equal to $O\left(\operatorname{polylog}\left(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})}\right)\right)$.
@@
Given a suitable approximation for the gap and the midpoint, we next use Lemma B.1 to prove forward error bounds between the true $k$-spectral projector of $\boldsymbol{\mathrm{M}}=\boldsymbol{\mathrm{\Phi}}^{\top}\boldsymbol{\mathrm{Q}}^{\top}\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{X}}^{\top}\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}$ and the one of $\widetilde{\boldsymbol{\mathrm{M}}}=\boldsymbol{\mathrm{M}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}$. From Lemma B.1, if we set $\mu=\widetilde{\mu}_{k}$, we have that $\|\operatorname{sgn}(\boldsymbol{\mathrm{M}}-\widetilde{\mu}_{k})-\operatorname{sgn}(\widetilde{\boldsymbol{\mathrm{M}}}-\widetilde{\mu}_{k})\|\leq\epsilon_{\mathsf{SGN}}$ if $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|\leq\epsilon_{\mathsf{SGN}}\tfrac{|\lambda_{\min}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{M}})|^{2}\pi}{128}=\epsilon_{\mathsf{SGN}}\Theta(\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})^{2})$. The same holds for the spectral projectors, i.e. $\|\boldsymbol{\mathrm{\Pi}}_{k}(\boldsymbol{\mathrm{M}})-\boldsymbol{\mathrm{\Pi}}_{k}(\widetilde{\boldsymbol{\mathrm{M}}})\|\leq\epsilon_{\mathsf{SGN}}$. The bound for $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|$ is achieved by setting $\textbf{{u}}\leq\frac{\epsilon_{\mathsf{SGN}}\widetilde{\operatorname{gap}}_{k}^{2}}{\operatorname{poly}(mnqk)\widetilde{\kappa}}$, which translates to $O\left(\log(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{SGN}}\widetilde{\operatorname{gap}}_{k}})\right)$ bits.
@@
We finally return $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}=\mathsf{MM}(\widetilde{\boldsymbol{\mathrm{Q}}},\widetilde{\boldsymbol{\mathrm{U}}}_{k})$. It holds that where $\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}-\widetilde{\boldsymbol{\mathrm{Q}}}\widetilde{\boldsymbol{\mathrm{U}}}_{k}\|\leq O(\operatorname{poly}(mqn)\textbf{{u}})$. We can then write
@@
where $\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}\boldsymbol{\mathrm{\Pi}}_{k}(\boldsymbol{\mathrm{M}})\boldsymbol{\mathrm{\Phi}}^{\top}\boldsymbol{\mathrm{Q}}^{\top}=\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}$ is equivalent to the true, exact arithmetic $\boldsymbol{\mathrm{Z}}_{k}$ of Algorithm 8. The error matrices satisfy $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Z}}}\|\leq O(\operatorname{poly}(mqn)\textbf{{u}})$, $\|\widetilde{\boldsymbol{\mathrm{Q}}}\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Pi}}}\widetilde{\boldsymbol{\mathrm{Q}}}^{\top}\|\leq O(\epsilon_{\mathsf{SGN}}+\epsilon_{\mathsf{PCA}})$, and $\|\boldsymbol{\mathrm{E}}^{\prime}\|\leq O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|)\in O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|)\in O(\epsilon_{\mathsf{SGN}}\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})^{2})$.
@@
Corollary G.1.
Using the Steps 1-6 as a floating point implementation of Algorithm 8, we can compute with high probability a matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\in\mathbb{R}^{m\times k}$ that satisfies
$\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\|\leq O(\epsilon_{\mathsf{PCA}}),$
where $\boldsymbol{\mathrm{Z}}_{k}$ is as in Algorithm 8. The total cost is at most
$O\left(qmnk^{\omega-2}\operatorname{polylog}(\tfrac{qmnk\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision.
@@
We first compute $\widetilde{\kappa}\in[\kappa(\boldsymbol{\mathrm{K}}),k^{c}\kappa(\boldsymbol{\mathrm{K}})]$ in Steps 1 and 2. This costs
$O(qmnk^{\omega-2}\operatorname{polylog}(qmnk\kappa(\boldsymbol{\mathrm{K}})))$
operations using $O(\operatorname{polylog}(qmnk\kappa(\boldsymbol{\mathrm{K}})))$ bits.
@@
$O\left(qmnk^{\omega-2}\log(1/\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})+(qk)^{\omega}\operatorname{polylog}(\tfrac{qk}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
@@
$O\left(\operatorname{polylog}(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
@@
In the last Step 6, assuming that $\epsilon_{\mathsf{SGN}}=\epsilon_{\mathsf{PCA}}$, we need to run Steps 1, 3, and 4, using $O\left(\operatorname{polylog}(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits. The number of arithmetic operations is not affected, only the bits. Thereafter the costs of $\mathsf{DEFLATE},$ $\mathsf{PURIFY}$, and $\mathsf{MM}$ are negligible compared to the costs of Steps 1, 3, and 4. The final matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}$ satisfies
$\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\|\leq O(\epsilon_{\mathsf{PCA}}),$
@@
$O\left(qmnk^{\omega-2}\operatorname{polylog}(\tfrac{qmnk\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision. ∎

### v4

@@ G.3 Bit complexity analysis of Block-Krylov PCA
In particular, we start with some desired bound $\epsilon^{\prime}_{0}\in(0,1)$ for $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|$ by setting the number of bits to $O(\log(\tfrac{qm\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}_{0}}))$. At each iteration we repeat Steps 1, 3, and 4 (we do not need to compute again the condition number of $\boldsymbol{\mathrm{K}}$ in Step 2). After at most $m=O(\log(1/\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})))$ iterations, the error $\epsilon^{\prime}_{m}$ satisfies $\epsilon^{\prime}_{m}=\Theta(\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}}))$, and we obtain two quantities $\widetilde{\operatorname{gap}}_{k}\in(1\pm\tfrac{1}{8})\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})$ and $\widetilde{\mu}_{k}\in\mu_{k}(\boldsymbol{\mathrm{M}})+\tfrac{1}{8}\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})$ with high probability. In each iteration we need to construct $\widetilde{\boldsymbol{\mathrm{M}}}$ using Steps 1, 3, and 4 with the specified number of bits.
Step 1 costs $O(qT_{\mathsf{MMX}}(k))$ operations. Step 3 requires $O(m(qk)^{\omega-1})$ operations. Step 4 executes $O(T_{\mathsf{MMX}}(qk)+m(qk)^{\omega-1})$ operations. The result is the $qk\times qk$ matrix $\widetilde{\boldsymbol{\mathrm{M}}}$ on which we call $\mathsf{COUNT}$ on each iteration. From Lemma E.1, assuming that we have regularized $\widetilde{\boldsymbol{\mathrm{M}}}$ appropriately using $\mathsf{REGULARIZE}$, $\mathsf{COUNT}$ costs at most $O((qk)^{\omega}\operatorname{polylog}(qk/\epsilon^{\prime}_{m}))=O((qk)^{\omega}\operatorname{polylog}(qk/\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})))$. The maximum number of bits is in the last iteration, which is equal to $O\left(\operatorname{polylog}\left(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})}\right)\right)$.
@@
Given a suitable approximation for the gap and the midpoint, we next use Lemma B.1 to prove forward error bounds between the true $k$-spectral projector of $\boldsymbol{\mathrm{M}}=\boldsymbol{\mathrm{\Phi}}^{\top}\boldsymbol{\mathrm{Q}}^{\top}\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{X}}^{\top}\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}$ and the one of $\widetilde{\boldsymbol{\mathrm{M}}}=\boldsymbol{\mathrm{M}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}$. From Lemma B.1, if we set $\mu=\widetilde{\mu}_{k}$, we have that $\|\operatorname{sgn}(\boldsymbol{\mathrm{M}}-\widetilde{\mu}_{k})-\operatorname{sgn}(\widetilde{\boldsymbol{\mathrm{M}}}-\widetilde{\mu}_{k})\|\leq\epsilon_{\mathsf{SGN}}$ if $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|\leq\epsilon_{\mathsf{SGN}}\tfrac{|\lambda_{\min}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{M}})|^{2}\pi}{128}=\epsilon_{\mathsf{SGN}}\Theta(\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})^{2})$. The same holds for the spectral projectors, i.e. $\|\boldsymbol{\mathrm{\Pi}}_{k}(\boldsymbol{\mathrm{M}})-\boldsymbol{\mathrm{\Pi}}_{k}(\widetilde{\boldsymbol{\mathrm{M}}})\|\leq\epsilon_{\mathsf{SGN}}$. The bound for $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|$ is achieved by setting $\textbf{{u}}\leq\frac{\epsilon_{\mathsf{SGN}}\widetilde{\operatorname{gap}}_{k}^{2}}{\operatorname{poly}(mq)\widetilde{\kappa}}$, which translates to $O\left(\log(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{SGN}}\widetilde{\operatorname{gap}}_{k}})\right)$ bits.
@@
We finally return $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}=\mathsf{MM}(\widetilde{\boldsymbol{\mathrm{Q}}},\widetilde{\boldsymbol{\mathrm{U}}}_{k})$. It holds that where $\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}-\widetilde{\boldsymbol{\mathrm{Q}}}\widetilde{\boldsymbol{\mathrm{U}}}_{k}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(mq))$. We can then write
@@
where $\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}\boldsymbol{\mathrm{\Pi}}_{k}(\boldsymbol{\mathrm{M}})\boldsymbol{\mathrm{\Phi}}^{\top}\boldsymbol{\mathrm{Q}}^{\top}=\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}$ is equivalent to the true, exact arithmetic $\boldsymbol{\mathrm{Z}}_{k}$ of Algorithm 8. The error matrices satisfy $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Z}}}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(mq))$, $\|\widetilde{\boldsymbol{\mathrm{Q}}}\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Pi}}}\widetilde{\boldsymbol{\mathrm{Q}}}^{\top}\|\leq O(\epsilon_{\mathsf{SGN}}+\epsilon_{\mathsf{PCA}})$, and $\|\boldsymbol{\mathrm{E}}^{\prime}\|\leq O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|)\in O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|)\in O(\epsilon_{\mathsf{SGN}}\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})^{2})$.
@@
Theorem G.2 (Restatement of Theorem 4.2).
Let $\boldsymbol{\mathrm{X}}$ be a data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, $\|\boldsymbol{\mathrm{X}}\|\leq 1$, $k\in[n]$ a target rank, $\epsilon_{\mathsf{PCA}}\in(0,1)$ an accuracy parameter, and $q=\Theta\left(\frac{\log(n)}{\sqrt{\epsilon_{\mathsf{PCA}}}}\right)$. Let $T_{\mathsf{MMX}}(k)$ denote the complexity to stably multiply $\boldsymbol{\mathrm{X}}$ or $\boldsymbol{\mathrm{X}}^{\top}$ with a dense matrix with $k$ columns from the right (see Def. G.1). Using the Steps 1-6 that are detailed in Appendix G.3 as a floating point implementation of Algorithm 8, we can compute a matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\in\mathbb{R}^{m\times k}$ that satisfies
$\left\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\right\|\leq O(\epsilon_{\mathsf{PCA}}),$
with high probability, where $\boldsymbol{\mathrm{Z}}_{k}$ is an approximate basis for the top-$k$ principal components of $\boldsymbol{\mathrm{X}}$, returned by Algorithm 8 in exact arithmetic. The total cost is at most
$O\left(qT_{\mathsf{MMX}}(k)\log(\tfrac{\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+m(qk)^{\omega-1}\log(\tfrac{1}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+(qk)^{\omega}\operatorname{polylog}(\tfrac{qk}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision. $\boldsymbol{\mathrm{K}},\boldsymbol{\mathrm{M}}$ are the same as in Alg. 8. The only parameters that are initially required are $k$ and $\epsilon_{\mathsf{PCA}}$.
@@
We first compute $\widetilde{\kappa}\in\Theta(\kappa(\boldsymbol{\mathrm{K}}))$ in Steps 1 and 2. This costs
$O\left(q\log(\kappa(\boldsymbol{\mathrm{K}}))T_{\mathsf{MMX}}(k)\right)$
floating point operations using $O(\operatorname{polylog}(qm\kappa(\boldsymbol{\mathrm{K}})))$ bits.
@@
$O\left(\left(qT_{\mathsf{MMX}}(k)+m(qk)^{\omega-1}\right)\log(\tfrac{1}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+(qk)^{\omega}\operatorname{polylog}(\tfrac{qk}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
@@
$O\left(\operatorname{polylog}(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
@@
In the last Step 6, assuming that $\epsilon_{\mathsf{SGN}}=\epsilon_{\mathsf{PCA}}$, we need to run Steps 1, 3, and 4, using $O\left(\operatorname{polylog}(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits. The number of arithmetic operations does not exceed the one from the previous step. Thereafter the costs of $\mathsf{DEFLATE},$ $\mathsf{PURIFY}$, and $\mathsf{MM}$ are negligible compared to the costs of Steps 1, 3, and 4. The final matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}$ satisfies
$\left\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\right\|\leq O(\epsilon_{\mathsf{PCA}}),$
@@
$O\left(qT_{\mathsf{MMX}}(k)\log(\tfrac{\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+m(qk)^{\omega-1}\log(\tfrac{1}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+(qk)^{\omega}\operatorname{polylog}(\tfrac{qk}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision. ∎